Please help me with this algebra question. Thank you so much in advance.
List (a) the coordinates of any y intercept
(b) the coordinates of any x intercept Do not graph.
10x - 3y = 60

Found 2 solutions by stanbon, algebrahouse.com:
Answer by stanbon(75887)

10x - 3y = 60
----
List
(a) the coordinates of any y intercept = ?
Let x = 0, then -3y=60 and y = -20
------------------------------------------
(b) the coordinates of any x intercept Do not graph.
Let y = 0, then 10x = 60 and x = 6
==========================================
Cheers,
Stan H.
===========

Answer by algebrahouse.com(1659)

10x - 3y = 60

To find the x-intercept (where the line crosses the x-axis), put 0 in for y:
10x - 3y = 60
10x - 3(0) = 60 {substituted 0 in for y}
10x = 60 {multiplied -3 by 0}
x = 6 {divided both sides by 10}
x-intercept is 6
coordinates of x-intercept are (6,0)

To find the y-intercept (where the line crosses the y-axis), put 0 in for x:
10x - 3y = 60
10(0) - 3y = 60 {substituted 0 in for x}
-3y = 60 {multiplied 10 by 0}
y = -20 {divided both sides by -3}
y-intercept is -20
coordinates of y-intercept are (-20,0)

For more help from me, visit: www.algebrahouse.com